OpenStudy (anonymous):

How would I go about this? Two ordered pairs on the hill are given as (6, 3) and (8, 7). Part 1: Find the slope of the hill. Slope (m) = change in y over change in x. Part 2: Now find the equation using the point slope formula. Point Slope Formula: y – y1 = m (x – x1) Part 3: Give an example of an equation for a line parallel to your answer from Part 2. Part 4: Give an example of an equation for a line perpendicular to your answer from Part 2.

OpenStudy (anonymous):

they practically tell you how to do it. For part 1, slope m is change in y over change in x. so subtract one y value from the other to get the change in y, and repeat for the x values. (6, 3) and (8, 7) y2 - y1 = 7 - 3 = 4 x2 - x1 = 8 - 6 = 2 slope = (y2 - y1) / (x2 - x1)

OpenStudy (anonymous):

So it would be 7-3/8-6?

OpenStudy (anonymous):

& that would be 2 correct?

OpenStudy (anonymous):

that is correct, slope is 2. You use that slope, m = 2, and either one of those points to work Part 2.

OpenStudy (anonymous):

How would I write the second part?

OpenStudy (anonymous):

They give you the formula again: Point Slope Formula: y – y1 = m (x – x1) So use m = 2 and substitute that in as your m. And pick one of those two points in the problem as (x1, y1)... let's use (6,3) as (x1, y1) Point Slope Formula: y – y1 = m (x – x1) y - 3 = 2 (x - 6)

OpenStudy (anonymous):

y = 2x - 9?

OpenStudy (anonymous):

Give an example written in slope intercept form (y = mx + b) of a line parallel to your solution in Part 2. So how would I write y = 2x - 9 of a line parallel to this? I know parallel lines always have the same slope..

OpenStudy (anonymous):

sorry, I had to step away. Yes, you're right... a parallel line has same slope, so your example line will have slope 2 also. It just has a different y-intercept. You can pick anything you want, except it needs to be different from the -9 in the Part 2... y = 2x - 9 y = 2x + 150 y = 2x - 3 these are just a couple examples...

OpenStudy (anonymous):

thank you :))